Determine the area under the standard normal curve that lies between ​

(a) Upper Z equals negative 0.45 and Upper Z equals 0.45​,

​(b) Upper Z equals negative 1.48 and Upper Z equals 0​, and

​(c) Upper Z equals negative 1.65 and Upper Z equals 2.23.
Answer #1

Using standard normal table,

(a)

P(-0.45 < z < 0.45)

= P(z < 0.45) - P(z < -0.45)

= 0.3473

Area = 0.3473

(b)

P(-1.48 < z < 0)

= P(z < 0) - P(z < -1.48)

= 0.4306

Area = 0.4306

(c)

P(-1.65 < z < 2.23)

= P(z < 2.23) - P(z < -1.65)

= 0.9377

Area = 0.9377